OpenStudy (anonymous):

-4y=-5x-18 linear equation medal will be given

OpenStudy (johnweldon1993):

So what are we doing? putting this into y = mx + b form?

OpenStudy (anonymous):

yeah so i have to have 3 solutions and all need a line

OpenStudy (anonymous):

on a graph and i cant graph a fraction or decmail it has to be a whole number

OpenStudy (johnweldon1993):

Oh okay...well you know that this is a linear equation (straight line) then when you have your 3 points...you can connect them...and hmm whole numbers...okay gimme 1 second :)

OpenStudy (johnweldon1993):

Okay 1 point...what does y equal when x = -2? 2nd point...what does y equal when x = 2?

OpenStudy (anonymous):

point one 7 point two 4

OpenStudy (johnweldon1993):

-4y=-5x-18 Point 1 (x = -2) -4y=-5(-2) -18 -4y = 10 - 18 -4y = -8 y = 2 so 1 point will be (-2 , 2) Point 2 (x = 2) -4y=-5(2) -18 -4y = -10 - 18 -4y = -28 y = 7 so another point will be (2 , 7) And now Point 3 (x = -6) -4y=-5(-6)-18 -4y = 30 - 18 -4y = 12 y = -3 so your third point will be (-6 , -3) Just plot those and then draw your straight line :)

OpenStudy (anonymous):

o gee thanks soo much :)

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

can u help meh with one that trips me up all the time??

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (anonymous):

2y=3x+11

OpenStudy (johnweldon1993):

same thing? find 3 points and graph?

OpenStudy (anonymous):

yuppers :)

OpenStudy (johnweldon1993):

Alright....so what does 'y' equal when x = 1? what does 'y' equal when x = 3? and what does 'y' equal when x = 5?

OpenStudy (anonymous):

point 1 is 7 point 2 is 10 point 3 is 13 ... but my graph doesnt go over ten so will it have to be a negitive??

OpenStudy (johnweldon1993):

Okay no problem...so keep the first 2 points and answer what does 'y' equal when x = -1?

OpenStudy (anonymous):

4

OpenStudy (johnweldon1993):

Perfect! So your 3 points are (-1 , 4) (1 , 7) and (3 , 10) So plot those and draw your line :)

OpenStudy (anonymous):

yay :) omgosh thanks sooo much

OpenStudy (johnweldon1993):

Haha see you already knew how to do it :)

OpenStudy (anonymous):

haha all i need was a little bit of guidence

OpenStudy (johnweldon1993):

And now you are perfect! :) Good job you :)

OpenStudy (anonymous):

nobodys perfect :) hahah and thanks :)

OpenStudy (johnweldon1993):

I meant with these types of questions :P and anytime :)

OpenStudy (anonymous):

oh right hahah i new that :P